We are CLOSED

Sensor Status Last Update
Open Switch Flipped OFF Oct 18, 2017 9:55 AM PDT
Front Door Closed Oct 18, 2017 4:59 PM PDT
Main Door Closed Oct 17, 2017 6:42 PM PDT
Office Motion Moving Oct 18, 2017 5:00 PM PDT
Shop Motion Moving Oct 18, 2017 5:00 PM PDT
Temperature 33.00 °C/ 91.40 °F Oct 18, 2017 5:04 PM PDT
Page Loaded at Oct 18, 2017 5:04 PM PDT